Carl has $135 in the bank and plans to save $5 per week. Jean has $90 in the bank and plans to save $10 per week. How many weeks will it be before Jean has at least as much in the bank as Carl?

Answer by rfer(16322)

135+5x=90+10x
45=5x
x=45/5
x=9 weeks